A fully charged 6.7 uF capacitor is connected in series with a 1.7×10^5 Ohm resistor.

A. What percentage of the original charge is left on the capacitor after 1.7 s of discharging?
Answer #1

Q = Qo (e-t/RC) t = 1.7 , R = 1.7*10^5 , C = 6.7 *10^-6

on putting we get

Q/Qo = 0.2248

percentage = 22.48 percent os charge is still left